OpenStudy (anonymous):

: Please help me solve this word problem I don't understand. Show an equation and a solution for the problem. Arnold needs a 25% solution of nitric acid. He has 20 ml of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution?

OpenStudy (mertsj):

.3(20)+.15x=.25(20+x)

OpenStudy (mertsj):

30(20)+15x=25(20+x)

OpenStudy (mertsj):

600+15x=500+25x

OpenStudy (mertsj):

100=10x

OpenStudy (mertsj):

10=x

OpenStudy (anonymous):

Thank you :D

OpenStudy (mertsj):

yw